OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Explain how you will show all of the solutions that satisfy this equation.Determine three different solutions for this equation. Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

Oh ok so I just substitute x=1 x=2 and x=3 and find all the solutions?

OpenStudy (anonymous):

In order to show all solutions you can either plot this equation or assume x = c, c is a real number and express y in terms of it: y = 9*c + 2, for any c let's find 3 solutions, set x = 0, we have: y = 2 set x =1: y = 9*1+2 = 11 set x = 2: y = 9*2+2 = 20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, in to find a equation that makes this system inconsistent, we rewrite your equation like: 9x - y = -2 y = 9x + 2 so it's slope is 9, a system is inconsistent if this two equations are parallel not coincident, so we must find a equation that has the same slope: so... y = 9x + h for any h, but 2

OpenStudy (anonymous):

you can choose any value for h, but not 2

OpenStudy (anonymous):

so, let's say h = 0 then: 9x - y = -2 9x - y = 0

OpenStudy (anonymous):

it's inconsistent

OpenStudy (nirmalnema):

condition for in consistence eq..is that two values have same x and y co ordinate but different constant

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

Thanks guys :) sorry for late response page reloaded lol